# Quantum many-body scars with unconventional superconducting pairing symmetries via multi-body interactions

### v2

@@ Abstract
We present a systematic framework to construct model Hamiltonians that have unconventional superconducting pairing states as exact energy eigenstates, by incorporating multi-body interactions (i.e., interactions among more than two particles). The multi-body interactions are introduced in a form of the local density-density coupling in such a way that any pair configuration in real space has a constant interaction energy by cancelling the two-body and multi-body interactions. Our approach is applicable to both spinless and spinful models in any spatial dimensions and on any bipartite lattices, facilitating an exhaustive extension of Yang’s $s$-wave $\eta$-pairing state to various other unconventional pairing symmetries ($p$-wave, $d$-wave, $f$-wave, etc.). Particularly, the constructed eigenstates have off-site pairs with finite center-of-mass momentum, which leads to superconducting states with either even-parity and spin-triplet or odd-parity and spin-singlet symmetry. We verify that the two-dimensional spinful Hubbard model on a square lattice with the multi-body interactions has the spin-triplet $d$-wave pairing state as an energy eigenstate, which can be regarded as a quantum many-body scar state as evidenced from the numerical analysis of the pair correlation function, entanglement entropy, and level statistics. We also discuss other examples, including spin-triplet $f$-wave pairing states on a honeycomb lattice and spin-singlet $p$-wave pairing states in a one dimensional chain. These findings open up the possibility of realizing nonequilibrium unconventional superconductivity in a long-lived manner protected against thermalization.
@@ I Introduction
Out-of-equilibrium superconductivity offers a potential to realize exotic quantum many-body states that are otherwise impossible to be found in thermal equilibrium. Notable candidates for those states include photoinduced superconducting-like states that have been experimentally observed even above the equilibrium transition temperature [1, 2, 3, 4, 5] (see also [6, 7]). Another example is the Floquet engineering [8, 9, 10] (i.e., quantum states induced by a time-periodic drive), which has endowed superconductors with topological properties [11, 12, 13, 14, 15, 16]. A variety of transient drives have also been employed to induce superconductivity with unconventional pairing states—such as bulk odd-frequency pairings [17, 18] and finite-momentum pairing states [19, 20]—which are challenging to be realized in equilibrium.
@@
Although numerous intriguing experimental and theoretical progress has been made, there still remains a problem of how such a nonthermal state can be made robust against thermalization. To generate nonequilibrium states, one necessarily injects a finite amount of energy to the system, which turns into heat in a general quantum many-body system. This results in the suppression (or vanishing) of superconducting correlations within a finite timescale. In order to overcome this difficulty, it is crucial to seek for a certain mechanism that protects nonequilibrium superconducting states from thermalization even in the presence of many-body interactions.
@@
As a promising mechanism to maintain superconducting correlations for a sufficiently long time at highly excited states, there have been proposed quantum many-body scar (QMBS) states [21, 22, 23], which are exceptional energy eigenstates that do not have thermal properties in a nonintegrable system. According to the eigenstate thermalization hypothesis (ETH) [24, 25, 26, 27], all excited energy eigenstates in a nonintegrable model are expected to be indistinguishable from thermal states as long as one refers to few-body observables. QMBS states are considered to be an exception to the ETH, as characterized by long-lived nonthermal dynamics in nonintegrable systems [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Several methods to construct a model Hamiltonian that accommodates a target state as a QMBS state have been proposed, such as the (restricted) spectrum-generating algebra [30], the embedding formalism [22], and the symmetry-based formalism [49].
@@
A primary example of a nonthermal energy eigenstate in a nonintegrable system that supports superconducting correlations is the so-called $\eta$-pairing state, which is an exact energy eigenstate of the Hubbard model on a $d$-dimensional square lattice, as shown by C. N. Yang [50]. Yang’s $\eta$-pairing state corresponds to a condensate of doublons carrying finite center-of-mass momentum, and exhibits long-range superconducting correlations with the $s$-wave and spin-singlet pairing symmetry. If one classifies the Hilbert space in terms of symmetry sectors, the $\eta$-pairing state is the only eigenstate in the corresponding sector [51, 52]. In this respect, strictly speaking, the $\eta$-pairing state in the Hubbard model is not considered to be a QMBS state, while there have been several proposals to modify the Hamiltonian to make it a true QMBS state [49, 53, 54, 55, 56, 57, 58, 59]. Similar $\eta$-pairing states have been recently proposed to be realized by photo-irradiation [19], and have attracted interest as a mechanism to induce superconductivity by light.
@@
Figure 1: (a), (b) Examples of spatial configurations of the spinless nearest-neighbor pairing state in a one-dimensional lattice. (a) When two pairs (blue clouds) are separated to each other, the two-body interaction $V$ only acts within each pair. (b) When two pairs come next to each other, there is an additional energy increment $V$, which can be cancelled by the three-body interaction $W$ (if one chooses $V+2W=0$). (c) Schematic picture of unconventional superconducting pairs of spinful fermions with $d$-wave (left panel) and $f$-wave (right) pairing symmetries (indicated by blue and red clouds) on the square and honeycomb lattices, respectively. Two pairs feel an energy cost $U$ due to the two-body interaction when they overlap with each other.
@@
To overcome this difficulty, we engineer multi-body interactions (i.e., interactions among more than two particles) to cancel the energy increment of pairs due to the two-body interaction. For example, in a one-dimensional spinless fermion model we can introduce a three-body interaction to cancel the nearest-neighbor two-body interaction acting on the pairs [Fig. 1(b)]. The cancellation works for arbitrary pair configurations, no matter how many pairs are distributed on the lattice. In this way, the spinless pairing state can be made an exact eigenstate of the one-dimensional Hubbard model with the three-body interaction. Previously, it has been proposed that the spinless pairing state becomes an exact eigenstate in a model with the three-body interaction and/or a density-dependent hopping [62, 64, 63], and a related model has been studied in the context of disordered quantum many-body scarred spin systems [62].
The advantage of our approach is that it can be systematically generalized to other pairing symmetries including spin degrees of freedom in higher dimensions [see Fig. 1(c)]. To this end, we introduce an extended number operator [$\mathbb{n}_{i\sigma}$ in Eq. (25)] that signals whether spin-$\sigma$ particles exist or not on lattice sites next to $i$ site. Using this operator, we can systematically cancel the energy increment induced by the formation of doublons when off-site pairs overlap with each other, and obtain a model Hamiltonian with multi-body interactions that has unconventional pairing states as exact eigenstates. We note that the derived interaction is strictly local, and takes a form of the density-density coupling between $M$ particles ($M\leq M_{\mathrm{max}}$ with $M_{\mathrm{max}}$ being finite and system-size independent).
@@
We apply our construction to the spinful Hubbard model with multi-body interactions on the two-dimensional square lattice, and show that the model has an exact eigenstate with the spin-triplet $d$-wave pairing symmetry [the left panel of Fig. 1(c)]. We numerically evidence that the model is nonintegrable from the analysis of energy level statistics. We also find that the $d$-wave pairing state shows the subvolume law of entanglement entropy and the off-diagonal long-range order of superconductivity, which indicates that the $d$-wave pairing state is indeed a QMBS state. Other examples are also discussed, including spin-triplet $f$-wave pairing states on a honeycomb lattice [the right panel of Fig. 1(c)] and spin-singlet $p$-wave pairing states in a one dimensional chain.
The rest of this paper is organized as follows. In Sec. II, we introduce a generalized $\eta$-pairing operator that creates off-site pairs with finite center-of-mass momentum, and classify its symmetry in terms of the parity, spin exchange, and sublattice exchange. In Sec. III, we demonstrate the idea of cancelling two-body interactions with multi-body ones to make the off-site pairing states exact eigenstates of a Hamiltonian. We will mainly focus on the case of the spin-triplet $d$-wave pairing symmetry on the two-dimensional square lattice as an example for explanation. We show that the unconventional pairing states become energy eigenstates of the multi-body interacting model. We numerically confirm that the unconventional pairing states are quantum many-body scar states based on the analysis of the pair correlation function, entanglement entropy (Sec. IV), and the level-spacing statistics (Sec. V). We will discuss generalizations to other pairing symmetries and other lattice structures, as well as extensions of the scarred models, in Sec. VI. Section VII contains a summary of the paper and an outlook for possible experimental realizations.
@@
Table 1: Symmetry classification of conventional and unconventional $\eta$-pairing states on a bipartite lattice, which are characterized by inversion ($\mathcal{P}=\pm 1$), spin exchange ($\mathcal{S}=\pm 1$), and sublattice exchange ($\varLambda=\pm 1$) symmetry. The product $\mathcal{P}\mathcal{S}\varLambda$ must satisfy $\mathcal{P}\mathcal{S}\varLambda=-1$. The sublattice symmetry $\varLambda$ distinguishes pairs on the same sublattice ($\varLambda=+1$) from those on the different sublattices ($\varLambda=-1$). In the case of the odd sublattice symmetry ($\varLambda=-1$), such as for nearest-neighbor (NN) pairings, a combination of the even-parity ($\mathcal{P}=+1$) and spin-triplet ($\mathcal{S}=+1$), or odd-parity ($\mathcal{P}=-1$) and spin-singlet ($\mathcal{S}=-1$) symmetry is allowed. In the case of the even sublattice symmetry ($\varLambda=+1$), such as for on-site and next-nearest-neighbor (NNN) pairings, a combination of the even-parity ($\mathcal{P}=+1$) and spin-singlet ($\mathcal{S}=-1$), or odd-parity ($\mathcal{P}=-1$) and spin-triplet ($\mathcal{S}=+1$) symmetry is allowed.
@@
Honeycomb $f$ wave Odd Triplet Even Sec. VI.2
@@ II Symmetry classification
Any pairing state must be anti-symmetric against the fermion exchange due to fermion’s anticommutation relation. For on-site pairs, for example, the anti-symmetry requires a combination of the spin-singlet and parity-even symmetry. Yang’s $\eta$-pairing state belongs to this class, being classified to the $s$-wave (parity even) and spin-singlet state. For off-site pairs spanned over different sublattices, on the other hand, there appears an additional minus sign under the fermion exchange, which comes from the alternating phase between different sublattices in Eq. (1). Exchanging the fermion operators within Eq. (1) and relabeling the dammy position variables, we obtain the following relation:
@@
one can fully specify the symmetry of the extended $\eta$-pairing operators in Eq. (1) under the restriction of $\mathcal{P}\mathcal{S}\varLambda=-1$ as imposed from the fermion anticommutation relation.
The symmetry classification of the on-site and off-site $\eta$-pairing states is summarised in Table 1 for various pairing symmetries discussed in this paper. For example, in a one-dimensional spinless fermion system, the nearest-neighbor $\eta$-pairing states, corresponding to the case of $\alpha=\pm 1$ with no internal degrees of freedom, are characterized by the $s$-wave ($\mathcal{P}=+1$), spin-triplet ($\mathcal{S}=+1$), and sublattice-odd ($\varLambda=-1$) pairing symmetry. This is in contrast to the spinless $p$-wave topological superconductivity in the Kitaev chain [74], which is even in the sublattice symmetry. These unconventional pairing symmetries originate from the staggered phase oscillation of the nonlocal pairing between different sublattices.
III Superconducting energy eigenstates with unconventional pairing symmetry
@@
In Sec. III.1, we take a glance at the simplest case of the spinless (or spin-triplet) $s$-wave pairing state in a one-dimensional system, which contains the essence of our idea utilizing multi-body interactions to construct the model Hamiltonians. We then discuss how the model construction can be generalized to other pairing symmetries. Here, we focus on a two-dimensional spinful model with multi-body interactions on the square lattice. In Sec. III.2, we will consider nearest-neighbor fermion pairs and introduce the $d$-wave spin-triplet $\eta$-pairing operator. In Sec. III.3, we see that the unconventional pairing states are not eigenstates of the two-body interacting model. In Sec. III.4, we introduce an extended number operator, which measures whether neighboring sites are occupied by particles. Using this operator, we can exactly and systematically cancel the energy increment between off-site pairs coming from the two-body interaction. We show that the spin-triplet $d$-wave pairing state becomes an exact eigenstate of the multi-body interacting model.

### v3

# Quantum many-body scars with unconventional superconducting pairing symmetries via multibody interactions

@@ Abstract
We present a systematic framework to construct model Hamiltonians that have unconventional superconducting pairing states as exact energy eigenstates, by incorporating multibody interactions (i.e., interactions among more than two particles). The multibody interactions are introduced in a form of the local density-density coupling in such a way that any pair configuration in real space has a constant interaction energy by canceling the two-body and multibody interactions. Our approach is applicable to both spinless and spinful models in any spatial dimensions and on any bipartite lattices, facilitating an exhaustive extension of Yang’s $s$-wave $\eta$-pairing state to various other unconventional pairing symmetries ($p$-wave, $d$-wave, $f$-wave, etc.). Particularly, the constructed eigenstates have off-site pairs with finite center-of-mass momentum, which leads to superconducting states with either even-parity and spin-triplet or odd-parity and spin-singlet symmetry. We verify that the two-dimensional spinful Hubbard model on a square lattice with the multibody interactions has the spin-triplet $d$-wave pairing state as an energy eigenstate, which can be regarded as a quantum many-body scar state as evidenced from the numerical analysis of the pair correlation function, entanglement entropy, and level statistics. We also discuss other examples, including spin-triplet $f$-wave pairing states on a honeycomb lattice and spin-singlet $p$-wave pairing states in a one-dimensional chain. These findings open up the possibility of realizing nonequilibrium unconventional superconductivity in a long-lived manner protected against thermalization.
@@ I Introduction
Out-of-equilibrium superconductivity offers a potential to realize exotic quantum many-body states that are otherwise challenging to observe in thermal equilibrium. Notable candidates for those states include photoinduced superconducting-like states that have been experimentally observed even above the equilibrium transition temperature [1, 2, 3, 4, 5] (see also [6, 7]). Another example is the Floquet engineering [8, 9, 10] (i.e., quantum states induced by a time-periodic drive), which has endowed superconductors with topological properties [11, 12, 13, 14, 15, 16]. A variety of transient drives have also been employed to induce superconductivity with unconventional pairing states—such as bulk odd-frequency pairings [17, 18] and finite-momentum pairing states [19, 20]—which are challenging to be realized in equilibrium.
@@
Although numerous intriguing experimental and theoretical progress has been made, there still remains a problem of how such a nonthermal state can be made robust against thermalization. To generate nonequilibrium states, one necessarily injects a finite amount of energy into the system, which typically leads to thermalization at elevated temperatures, thereby suppressing (or even erasing) superconducting correlations. In order to overcome this difficulty, it is crucial to seek for a certain mechanism that protects nonequilibrium superconducting states from thermalization even in the presence of many-body interactions.
@@
As a promising mechanism to maintain superconducting correlations for a sufficiently long time at highly excited states, there have been proposed quantum many-body scar (QMBS) states [21, 22, 23], which are exceptional energy eigenstates that do not have thermal properties in a nonintegrable system. According to the eigenstate thermalization hypothesis (ETH) [24, 25, 26, 27], all excited energy eigenstates in a nonintegrable model are expected to be indistinguishable from thermal states as long as one refers to few-body observables. QMBS states are considered to be an exception to the ETH, as characterized by long-lived nonthermal dynamics in nonintegrable systems [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Despite being exact energy eigenstates in a static Hamiltonian, QMBS states differ from equilibrium states: In equilibrium, if one takes the microcanonical ensemble, most of the eigenstates in the energy shell are thermal (according to the ETH). Even if the QMBS states are included in the shell, they are overwhelmed by those thermal states. Several methods have been proposed to construct a model Hamiltonian that accommodates a target state as a QMBS state, such as the (restricted) spectrum-generating algebra [30], the embedding formalism [22], and the symmetry-based formalism [49].
@@
A primary example of a nonthermal energy eigenstate in a nonintegrable system that supports superconducting correlations is the so-called $\eta$-pairing state, which is an exact energy eigenstate of the Hubbard model on a $d$-dimensional square lattice, as shown by C. N. Yang [50]. Yang’s $\eta$-pairing state corresponds to a condensate of doublons carrying finite center-of-mass momentum, and exhibits long-range superconducting correlations with the $s$-wave and spin-singlet pairing symmetry. If one classifies the Hilbert space in terms of symmetry sectors, the $\eta$-pairing state is the only eigenstate in the corresponding sector, characterized by the so-called $\eta$-$\mathrm{SU}(2)$ symmetry [51, 52]. In this respect, several proposals have been made to modify the Hamiltonian to break the $\eta$-$\mathrm{SU}(2)$ symmetry, thereby transforming the Yang’s $\eta$-pairing state into a QMBS state [49, 53, 54, 55, 56, 57, 58, 59]. Similar $\eta$-pairing states have recently been proposed to be realized by photoirradiation [19], and have attracted interest as a mechanism to induce superconductivity by light. Once the system is excited to the $\eta$-pairing states, the system in principle stays there forever (since they are the exact eigenstates). Moreover, in nonequilibrium situations one might expect to see rich superconducting states with unconventional pairing symmetries that are not typically found in equilibrium.
@@
Figure 1: (a), (b) Examples of spatial configurations of the spinless nearest-neighbor pairing state in a one-dimensional lattice. (a) When two pairs (blue clouds) are separated to each other, the two-body interaction $V$ only acts within each pair. (b) When two pairs come next to each other, there is an additional energy increment $V$, which can be canceled by the three-body interaction $W$ (if one chooses $V+2W=0$). (c) Schematic picture of unconventional superconducting pairs of spinful fermions with $d$-wave (left panel) and $f$-wave (right) pairing symmetries (indicated by blue and red clouds) on the square and honeycomb lattices, respectively. Two pairs feel an energy cost $U$ due to the two-body interaction when they overlap with each other.
@@
To overcome this difficulty, we engineer multibody interactions (i.e., interactions among more than two particles) to cancel the energy increment of pairs due to the two-body interaction. For example, in a one-dimensional spinless fermion model we can introduce a three-body interaction to cancel the nearest-neighbor two-body interaction acting on the pairs [Fig. 1(b)]. The cancellation works for arbitrary pair configurations, no matter how many pairs are distributed on the lattice. In this way, the spinless pairing state can be made an exact eigenstate of the one-dimensional Hubbard model with the three-body interaction. Previously, it has been proposed that the spinless pairing state becomes an exact eigenstate in a model with the three-body interaction and/or a density-dependent hopping [54, 62, 64, 63], and a related model has been studied in the context of disordered quantum many-body scarred spin systems [62].
The advantage of our approach is that it can be systematically generalized to other pairing symmetries including spin degrees of freedom in higher dimensions [see Fig. 1(c)]. To this end, we introduce an extended number operator [$\mathbb{n}_{i\sigma}$ in Eq. (25)] that signals whether spin-$\sigma$ particles exist or not on lattice sites next to $i$ site. Using this operator, we can systematically cancel the energy increment induced by the formation of doublons when off-site pairs overlap with each other, and obtain a model Hamiltonian with multibody interactions that has unconventional pairing states as exact eigenstates. We note that the derived interaction is strictly local, and takes a form of the density-density coupling between $M$ particles ($M\leq M_{\mathrm{max}}$ with $M_{\mathrm{max}}$ being finite and system-size independent).
@@
We apply our construction to the spinful Hubbard model with multibody interactions on the two-dimensional square lattice, and show that the model has an exact eigenstate with the spin-triplet $d$-wave pairing symmetry [the left panel of Fig. 1(c)]. We numerically evidence that the model is nonintegrable from the analysis of energy level statistics. We also find that the $d$-wave pairing state shows the subvolume law of entanglement entropy and the off-diagonal long-range order of superconductivity, which indicates that the $d$-wave pairing state is indeed a QMBS state. Other examples are also discussed, including spin-triplet $f$-wave pairing states on a honeycomb lattice [the right panel of Fig. 1(c)] and spin-singlet $p$-wave pairing states in a one dimensional chain.
The rest of this paper is organized as follows. In Sec. II, we introduce a generalized $\eta$-pairing operator that creates off-site pairs with finite center-of-mass momentum, and classify its symmetry in terms of the parity, spin exchange, and sublattice exchange. In Sec. III, we demonstrate the idea of cancelling two-body interactions with multibody ones to make the off-site pairing states exact eigenstates of a Hamiltonian. We will mainly focus on the case of the spin-triplet $d$-wave pairing symmetry on the two-dimensional square lattice as an example for explanation. We show that the unconventional pairing states become energy eigenstates of the multibody interacting model. We numerically confirm that the unconventional pairing states are quantum many-body scar states based on the analysis of the pair correlation function, entanglement entropy (Sec. IV), and the level-spacing statistics (Sec. V). We will discuss generalizations to other pairing symmetries and other lattice structures, as well as extensions of the scarred models and the degeneracy structure of those pairing states, in Sec. VI. Section VII contains a summary of the paper and an outlook for possible experimental realizations.
@@
Table 1: Symmetry classification of conventional and unconventional $\eta$-pairing states on a bipartite lattice, which are characterized by inversion [$\mathcal{P}=\pm 1$, Eq. (6)], spin exchange [$\mathcal{S}=\pm 1$, Eq. (7)], and sublattice exchange [$\varLambda=\pm 1$, Eq. (5b)] symmetry. The product $\mathcal{P}\mathcal{S}\varLambda$ must satisfy $\mathcal{P}\mathcal{S}\varLambda=-1$. The sublattice symmetry $\varLambda$ distinguishes pairs on the same sublattice ($\varLambda=+1$) from those on the different sublattices ($\varLambda=-1$). In the case of the odd sublattice symmetry ($\varLambda=-1$), such as for nearest-neighbor (NN) pairings, a combination of the even-parity ($\mathcal{P}=+1$) and spin-triplet ($\mathcal{S}=+1$), or odd-parity ($\mathcal{P}=-1$) and spin-singlet ($\mathcal{S}=-1$) symmetry is allowed. In the case of the even sublattice symmetry ($\varLambda=+1$), such as for on-site and next-nearest-neighbor (NNN) pairings, a combination of the even-parity ($\mathcal{P}=+1$) and spin-singlet ($\mathcal{S}=-1$), or odd-parity ($\mathcal{P}=-1$) and spin-triplet ($\mathcal{S}=+1$) symmetry is allowed. * indicates that the mirror transformation is considered instead of the parity transformation.
@@
Honeycomb $f$ wave Odd Triplet Even Sec. VI.2, ${}^{*}\!$irror symmetry
@@ II Symmetry classification
Any pairing state must be anti-symmetric against the fermion exchange due to fermion’s anticommutation relation. For on-site pairs, for example, the anti-symmetry requires a combination of the spin-singlet and parity-even symmetry. Yang’s $\eta$-pairing state belongs to this class, being classified to the $s$-wave (parity even) and spin-singlet state. For off-site pairs spanned over different sublattices, on the other hand, there appears an additional minus sign under the fermion exchange, which comes from the alternating phase between different sublattices in Eq. (1). Exchanging the fermion operators within Eq. (1) and relabeling the dammy position variables $\bm{r}_{i}+\bm{\alpha}$ as the new ones $\bm{r}^{\prime}_{i}$, we obtain the following relation:
@@
one can symmetrize the extended $\eta$-pairing operators (1) via appropriate linear combinations, thereby fully specifying their symmetries under the restriction of $\mathcal{P}\mathcal{S}\varLambda=-1$ as imposed from the fermion anticommutation relation.
The symmetry classification of the on-site and off-site $\eta$-pairing states is summarized in Table 1 for various pairing symmetries discussed in this paper. For example, in a one-dimensional spinless fermion system, the nearest-neighbor $\eta$-pairing states, corresponding to the case of $\alpha=\pm 1$ with no internal degrees of freedom, are characterized by the $s$-wave ($\mathcal{P}=+1$), spin-triplet ($\mathcal{S}=+1$), and sublattice-odd ($\varLambda=-1$) pairing symmetry. This is in contrast to the spinless $p$-wave topological superconductivity in the Kitaev chain [74], which is even in the sublattice symmetry. These unconventional pairing symmetries originate from the staggered phase oscillation of the nonlocal pairing between different sublattices.
We remark the case where the site-centered parity symmetry is absent. A representative example is the honeycomb lattice, where the site-symmetry group (i.e., the point group that fixes one lattice site) is $C_{3v}$ [75], which does not include the parity symmetry. In those cases, one can consider the mirror symmetry instead of the parity symmetry (Table 1). Actually, the irreducible representation corresponding to the $f$-wave symmetry belongs to $A_{2}$, which is odd under the mirror symmetry, being distinguished from other irreducible representations.
III Superconducting energy eigenstates with unconventional pairing symmetries
@@
In Sec. III.1, we take a glance at the simplest case of the spinless (or spin-triplet) $s$-wave pairing state in a one-dimensional system, which contains the essence of our idea utilizing multibody interactions to construct the model Hamiltonians. We then discuss how the model construction can be generalized to other pairing symmetries. Here, we focus on a two-dimensional spinful model with multibody interactions on the square lattice. In Sec. III.2, we will consider nearest-neighbor fermion pairs and introduce the $d$-wave spin-triplet $\eta$-pairing operator. In Sec. III.3, we see that the unconventional pairing states are not eigenstates of the two-body interacting model. In Sec. III.4, we introduce an extended number operator, which measures whether neighboring sites are occupied by particles. Using this operator, we can exactly and systematically cancel the energy increment between off-site pairs coming from the two-body interaction. We show that the spin-triplet $d$-wave pairing state becomes an exact eigenstate of the multibody interacting model.